OpenStudy (anonymous):

Two stores sell the same computer for the same original price. Store A advertises that the computer is on sale for 25% off the original price. Store B advertises that it is reducing the computer’s price by $180. When Brittany compares the sale prices of the computer in both stories, she concludes that the sale prices are equal. Which equation models this situation? Let p represent the computer’s original price. A. 0.25p = p – 180 B. 0.75p = p + 180 C. 0.25(p – 180) = p D. 0.75p = p − 180

OpenStudy (anonymous):

p - 25% of p is the same as p - 180 p - 0.25p = p -180 0.75p=p-180 (D)